# Group Representational Position Encoding

### v3

@@ 1 Introduction
    We propose GRAPE as a unified group‑theoretic view that subsumes multiplicative orthogonal rotations in $\operatorname{SO}(d)$ and additive unipotent (all eigenvalues equal to 11) mechanisms in general linear group $\mathrm{GL}$, recovering RoPE and ALiBi as exact special cases and proving FoX is an exact instance (Appendix C).
@@ 2.1 Preliminaries and Rank-2 Generator
Initialization. Write $\mathbf{A}\triangleq[\mathbf{a}\ \mathbf{b}]\in\mathbb{R}^{d\times 2}$ and $\mathbf{J}=\begin{pmatrix}0&-1\\[0.86108pt] 1&0\end{pmatrix}$ so that $\mathbf{L}=\mathbf{A}\mathbf{J}\mathbf{A}^{\top}$. For any $\mathbf{M}\in\mathrm{SL}(2)$ (the $2\times 2$ real matrices with determinant 11, see Table 3), $\mathbf{M}\mathbf{J}\mathbf{M}^{\top}=\mathbf{J}$ and thus $\mathbf{A}\mapsto\mathbf{A}\mathbf{M}$ leaves $\mathbf{L}$ invariant; for general $\mathbf{M}\in\mathrm{GL}(2)$ the group of invertible $2\times 2$ matrices), $\mathbf{L}$ scales by $\det(\mathbf{M})$. Therefore the oriented plane $\mathcal{U}=\mathrm{span}\{\mathbf{a},\mathbf{b}\}$ and the scalar $s=\sqrt{\alpha\beta-\gamma^{2}}$ determine the action. We fix a gauge at initialization by $\|\mathbf{a}\|=\|\mathbf{b}\|=1$ and $\mathbf{a}^{\top}\mathbf{b}=0$ (absorbing scale into $\omega$).
@@ 3.1 Multi‑Subspace GRAPE-M and RoPE as a Special Case
Spectral parameterization. Classical RoPE chooses $\theta_{i}$ on a log‑uniform grid across ii. In GRAPE, $\theta_{i}$ can be learned or shared/tied across heads or layers. The MS‑GRAPE view also allows replacing the coordinate selectors $\mathbf{U}_{i}$ by a learned orthogonal basis $\mathbf{B}\in\operatorname{SO}(d)$ so that $\mathbf{L}=\sum_{i}\theta_{i}\mathbf{B}\mathbf{U}_{i}\mathbf{J}\mathbf{U}_{i}^{\top}\mathbf{B}^{\top}$, preserving commutativity while learning subspaces.
@@ 4 Additive Group Representational Position Encoding
This section shows that additive positional mechanisms (absolute shifts of features and additive logit biases, including ALiBi (Press et al., 2021)) also admit a group-theoretic formulation. The key is a homogeneous lift to an augmented space and a one-parameter subgroup of the general linear group $\mathrm{GL}$ that acts by unipotent (all eigenvalues equal to 11) transformations. This yields an exact relative law and streaming/cache rules analogous to Section 2.5.
@@ 4.1 Homogeneous lift and a unipotent action
Canonical rank-1 generator in a (d+1)(d{+}1)-lift. A convenient concrete choice (used in Appendix J.3) is the rank-11 index-22 nilpotent
@@
Note that under the asymmetric lift $\widehat{\mathbf{q}}=[\mathbf{q};1;0]$, $\widehat{\mathbf{k}}=[\mathbf{k};0;1]$, the base inner product satisfies $\widehat{\mathbf{q}}_{i}^{\top}\widehat{\mathbf{k}}_{j}=\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$ (no extra constant), while the rank-11 update contributes exactly $(j{-}i)\omega(\lambda_{q}+\lambda_{k})$, which is non-positive for causal offsets $j\leq i$, which we denoted as GRAPE-A-QK. This derives a learnable, content-adaptive linear bias (slope) from first principles: it is the sum of query-dependent and key-dependent unipotent actions in the general linear group.
@@ 5 Path Integral Additive GRAPE
Definition (Path‑integral bias). Fix a head $h$ and per-head scale $\alpha_{h}>0$. For each time $u$, let $\mathbf{p}_{u,h}\in\mathbb{R}^{d}$ be a positional embedding obtained from token-local features (a linear projection followed by RMS normalization in our implementation). Let $\mathcal{J}$ be the canonical block-diagonal $90^{\circ}$ operator (Section 2.4), and define $\mathbf{R}_{\ell}:=\exp(\ell\,\mathcal{J})$ (a fixed commuting rotation). For a link function $g:\mathbb{R}\to(-\infty,0)$ that is monotone increasing and 11-Lipschitz [^2] Our experiments take $g(z)=\log(\text{Sigmoid}(z))$; then $g^{\prime}(z)=1-\text{Sigmoid}(z)\in(0,1)$, ensuring 11-Lipschitzness., define the edge potential
@@
Group‑theoretic formalization and path composition. Let $\mathbf{E}\in\mathbb{R}^{(d{+}2)\times(d{+}2)}$ be a fixed rank-11 nilpotent with $\mathbf{E}^{2}=\mathbf{0}$ (e.g., $\mathbf{E}=\mathbf{e}_{d+2}\mathbf{e}_{d+1}^{\top}$ as in Section 4.2). For each fixed endpoint $t$, define endpoint-indexed unipotent factors
@@
Spectral and stability. Each factor $\mathbf{H}_{h}^{(t)}(\ell)=\mathbf{I}-\psi_{h}(t,\ell)\mathbf{E}$ is unipotent with all eigenvalues 11 and at most two singular values deviating from 11; the full path product equals $\mathbf{I}-b_{h}(t,j)\mathbf{E}$ (Eq. (5.4)). As in Appendix J.3, the paired inverse-transpose used for scoring cancels multiplicative distortions and delivers exactly the additive bias $b_{h}(t,j)$; operator norms remain controlled linearly in $|b_{h}(t,j)|$.
@@ Appendix I Differentiation and Fast Application of Rank-2 Matrix Exponential
We now analyze the spectral properties of the additive lifts in $\mathrm{GL}$ introduced in Sections 4 and 5. The key structural fact is unipotency: all per-step factors are identity plus a rank-1 (or few-rank) nilpotent update of index 22.
@@
Hence, the minimal polynomial of $\mathbf{H}(s)$ is $(\lambda-1)^{2}$, and the Jordan form consists of size-22 Jordan blocks for the 11-eigenspace, with the number of nontrivial blocks equal to $\operatorname{rank}(\mathbf{A})$.
@@
Since $\mathbf{A}^{2}=\mathbf{0}$, $\exp(s\mathbf{A})=\mathbf{I}+s\mathbf{A}$ and $(\mathbf{H}(s)-\mathbf{I})^{2}=s^{2}\mathbf{A}^{2}=\mathbf{0}$. The characteristic polynomial is $(\lambda-1)^{D}$ for $\mathbf{H}(s)$, so all eigenvalues equal 11. The determinant equals the product of eigenvalues, hence 11; the spectral radius is therefore 11. ∎
@@
so the combined generator is also index-22 nilpotent and yields the same unipotent spectrum.
@@
Let $\mathbf{E}:=\mathbf{e}_{p}\,\mathbf{e}_{q}^{\top}$ with $p\neq q$ and define $\mathbf{H}(s):=\mathbf{I}+s\mathbf{E}\in\mathbb{R}^{D\times D}$. Then D2D-2 singular values equal 11, and the remaining two are
@@
Lemma J.6 (General operator-norm bounds for index-22 unipotents).
@@
These bounds are conservative but dimension-free. In the canonical rank-11 case of Lemma J.4 with $\|\mathbf{A}\|_{2}=1$, one has the sharper small-$|s|$ behavior $\sigma_{\max}(\mathbf{I}+s\mathbf{A})=1+\tfrac{|s|}{2}+O(s^{2})$ and $\sigma_{\min}(\mathbf{I}+s\mathbf{A})=1-\tfrac{|s|}{2}+O(s^{2})$.
@@
because $\mathbf{E}^{2}=\mathbf{0}$. Thus GRAPE-AP inherits the unipotent spectrum of Prop. J.3 with row-dependent $s=s_{h}(t,j)\leq 0$ (since $\psi_{h}\leq 0$ by construction). Its only two nontrivial singular values are exactly (J.1) with $s\mapsto s_{h}(t,j)$; the rest equal 11. Consequently,
@@
while the determinant remains 11 and eigenvalues are all 11. As in Additive GRAPE(GRAPE-A), the paired inverse-transpose used in the bilinear scoring removes any multiplicative anisotropy, leaving the bounded additive term $b_{h}(t,j)$ in Eq. (5.2).
@@
Thus $\mathbf{H}_{t}$ is norm nonexpansive (operator norm 11) but not norm-preserving unless $\beta_{t}\in\{0,2\}$. Singular values equal the absolute eigenvalues because $\mathbf{H}_{t}$ is symmetric; the component along $\mathbf{w}_{t}$ is scaled by $|1-\beta_{t}|<1$ for any $\beta_{t}\in(0,2)\setminus\{0,2\}$, and flips sign when $\beta_{t}>1$ (a design choice in PaTH to allow negative eigenvalues for state-tracking).
@@
Hence $\mathbf{P}_{j\to i}$ is (at best) nonexpansive, with a worst-case exponential lower bound on the smallest singular value governed by the path-length product of $|1-\beta_{s}|$. Whenever some $\beta_{s}$ is close to 11, $\mathbf{H}_{s}$ is nearly singular (and exactly singular if $\beta_{s}=1$), driving $\sigma_{\min}(\mathbf{P}_{j\to i})$ toward zero. Volume contraction is quantified by
@@
Implications for long-context modeling. Because the PaTH transport multiplies the Q/K bilinear by $\mathbf{P}_{j\to i}$, any persistent deviation of $\beta_{t}$ from $\{0,2\}$ yields cumulative energy loss along a moving one-dimensional subspace. This concentrates mass in progressively fewer directions and can flatten or attenuate long-range logits $\mathbf{q}_{i}^{\top}\mathbf{P}_{j\to i}\mathbf{k}_{j}$ as iji-j grows, unless additional renormalizations or forget-gates are introduced. In contrast, GRAPE-M maps lie in $\operatorname{SO}(d)$, so for both non‑contextual and contextual types, all singular values are 11; volumes and norms are preserved, and Lyapunov exponents are 0, avoiding contraction‑induced degradation of long‑range interactions.

### v4

@@ 1 Introduction
    We propose GRAPE as a unified group‑theoretic view that subsumes multiplicative orthogonal rotations in $\operatorname{SO}(d)$ and additive unipotent (all eigenvalues equal to $1$) mechanisms in general linear group $\mathrm{GL}$, recovering RoPE and ALiBi as exact special cases and proving FoX is an exact instance (Appendix C).
@@ 2.1 Preliminaries and Rank-2 Generator
Initialization. Write $\mathbf{A}\triangleq[\mathbf{a}\ \mathbf{b}]\in\mathbb{R}^{d\times 2}$ and $\mathbf{J}=\begin{pmatrix}0&-1\\[0.86108pt] 1&0\end{pmatrix}$ so that $\mathbf{L}=\mathbf{A}\mathbf{J}\mathbf{A}^{\top}$. For any $\mathbf{M}\in\mathrm{SL}(2)$ (the $2\times 2$ real matrices with determinant $1$, see Table 3), $\mathbf{M}\mathbf{J}\mathbf{M}^{\top}=\mathbf{J}$ and thus $\mathbf{A}\mapsto\mathbf{A}\mathbf{M}$ leaves $\mathbf{L}$ invariant; for general $\mathbf{M}\in\mathrm{GL}(2)$ the group of invertible $2\times 2$ matrices), $\mathbf{L}$ scales by $\det(\mathbf{M})$. Therefore the oriented plane $\mathcal{U}=\mathrm{span}\{\mathbf{a},\mathbf{b}\}$ and the scalar $s=\sqrt{\alpha\beta-\gamma^{2}}$ determine the action. We fix a gauge at initialization by $\|\mathbf{a}\|=\|\mathbf{b}\|=1$ and $\mathbf{a}^{\top}\mathbf{b}=0$ (absorbing scale into $\omega$).
@@ 3.1 Multi‑Subspace GRAPE-M and RoPE as a Special Case
Spectral parameterization. Classical RoPE chooses $\theta_{i}$ on a log‑uniform grid across $i$. In GRAPE, $\theta_{i}$ can be learned or shared/tied across heads or layers. The MS‑GRAPE view also allows replacing the coordinate selectors $\mathbf{U}_{i}$ by a learned orthogonal basis $\mathbf{B}\in\operatorname{SO}(d)$ so that $\mathbf{L}=\sum_{i}\theta_{i}\mathbf{B}\mathbf{U}_{i}\mathbf{J}\mathbf{U}_{i}^{\top}\mathbf{B}^{\top}$, preserving commutativity while learning subspaces.
@@ 4 Additive Group Representational Position Encoding
This section shows that additive positional mechanisms (absolute shifts of features and additive logit biases, including ALiBi (Press et al., 2021)) also admit a group-theoretic formulation. The key is a homogeneous lift to an augmented space and a one-parameter subgroup of the general linear group $\mathrm{GL}$ that acts by unipotent (all eigenvalues equal to $1$) transformations. This yields an exact relative law and streaming/cache rules analogous to Section 2.5.
@@ 4.1 Homogeneous lift and a unipotent action
Canonical rank-1 generator in a (d+1)(d{+}1)-lift. A convenient concrete choice (used in Appendix J.3) is the rank-$1$ index-22 nilpotent
@@
Note that under the asymmetric lift $\widehat{\mathbf{q}}=[\mathbf{q};1;0]$, $\widehat{\mathbf{k}}=[\mathbf{k};0;1]$, the base inner product satisfies $\widehat{\mathbf{q}}_{i}^{\top}\widehat{\mathbf{k}}_{j}=\mathbf{q}_{i}^{\top}\mathbf{k}_{j}$ (no extra constant), while the rank-$1$ update contributes exactly $(j{-}i)\omega(\lambda_{q}+\lambda_{k})$, which is non-positive for causal offsets $j\leq i$, which we denoted as GRAPE-A-QK. This derives a learnable, content-adaptive linear bias (slope) from first principles: it is the sum of query-dependent and key-dependent unipotent actions in the general linear group.
@@ 5 Path Integral Additive GRAPE
Definition (Path‑integral bias). Fix a head $h$ and per-head scale $\alpha_{h}>0$. For each time $u$, let $\mathbf{p}_{u,h}\in\mathbb{R}^{d}$ be a positional embedding obtained from token-local features (a linear projection followed by RMS normalization in our implementation). Let $\mathcal{J}$ be the canonical block-diagonal $90^{\circ}$ operator (Section 2.4), and define $\mathbf{R}_{\ell}:=\exp(\ell\,\mathcal{J})$ (a fixed commuting rotation). For a link function $g:\mathbb{R}\to(-\infty,0)$ that is monotone increasing and $1$-Lipschitz [^2] Our experiments take $g(z)=\log(\text{Sigmoid}(z))$; then $g^{\prime}(z)=1-\text{Sigmoid}(z)\in(0,1)$, ensuring $1$-Lipschitzness., define the edge potential
@@
Group‑theoretic formalization and path composition. Let $\mathbf{E}\in\mathbb{R}^{(d{+}2)\times(d{+}2)}$ be a fixed rank-$1$ nilpotent with $\mathbf{E}^{2}=\mathbf{0}$ (e.g., $\mathbf{E}=\mathbf{e}_{d+2}\mathbf{e}_{d+1}^{\top}$ as in Section 4.2). For each fixed endpoint $t$, define endpoint-indexed unipotent factors
@@
Spectral and stability. Each factor $\mathbf{H}_{h}^{(t)}(\ell)=\mathbf{I}-\psi_{h}(t,\ell)\mathbf{E}$ is unipotent with all eigenvalues $1$ and at most two singular values deviating from $1$; the full path product equals $\mathbf{I}-b_{h}(t,j)\mathbf{E}$ (Eq. (5.4)). As in Appendix J.3, the paired inverse-transpose used for scoring cancels multiplicative distortions and delivers exactly the additive bias $b_{h}(t,j)$; operator norms remain controlled linearly in $|b_{h}(t,j)|$.
@@ Appendix I Differentiation and Fast Application of Rank-2 Matrix Exponential
We now analyze the spectral properties of the additive lifts in $\mathrm{GL}$ introduced in Sections 4 and 5. The key structural fact is unipotency: all per-step factors are identity plus a rank-1 (or few-rank) nilpotent update of index $2$.
@@
Hence, the minimal polynomial of $\mathbf{H}(s)$ is $(\lambda-1)^{2}$, and the Jordan form consists of size-$2$ Jordan blocks for the $1$-eigenspace, with the number of nontrivial blocks equal to $\operatorname{rank}(\mathbf{A})$.
@@
Since $\mathbf{A}^{2}=\mathbf{0}$, $\exp(s\mathbf{A})=\mathbf{I}+s\mathbf{A}$ and $(\mathbf{H}(s)-\mathbf{I})^{2}=s^{2}\mathbf{A}^{2}=\mathbf{0}$. The characteristic polynomial is $(\lambda-1)^{D}$ for $\mathbf{H}(s)$, so all eigenvalues equal $1$. The determinant equals the product of eigenvalues, hence $1$; the spectral radius is therefore $1$. ∎
@@
so the combined generator is also index-$2$ nilpotent and yields the same unipotent spectrum.
@@
Let $\mathbf{E}:=\mathbf{e}_{p}\,\mathbf{e}_{q}^{\top}$ with $p\neq q$ and define $\mathbf{H}(s):=\mathbf{I}+s\mathbf{E}\in\mathbb{R}^{D\times D}$. Then D2D-2 singular values equal $1$, and the remaining two are
@@
Lemma J.6 (General operator-norm bounds for index-$2$ unipotents).
@@
These bounds are conservative but dimension-free. In the canonical rank-$1$ case of Lemma J.4 with $\|\mathbf{A}\|_{2}=1$, one has the sharper small-$|s|$ behavior $\sigma_{\max}(\mathbf{I}+s\mathbf{A})=1+\tfrac{|s|}{2}+O(s^{2})$ and $\sigma_{\min}(\mathbf{I}+s\mathbf{A})=1-\tfrac{|s|}{2}+O(s^{2})$.
@@
because $\mathbf{E}^{2}=\mathbf{0}$. Thus GRAPE-AP inherits the unipotent spectrum of Prop. J.3 with row-dependent $s=s_{h}(t,j)\leq 0$ (since $\psi_{h}\leq 0$ by construction). Its only two nontrivial singular values are exactly (J.1) with $s\mapsto s_{h}(t,j)$; the rest equal $1$. Consequently,
@@
while the determinant remains $1$ and eigenvalues are all $1$. As in Additive GRAPE(GRAPE-A), the paired inverse-transpose used in the bilinear scoring removes any multiplicative anisotropy, leaving the bounded additive term $b_{h}(t,j)$ in Eq. (5.2).
@@
Thus $\mathbf{H}_{t}$ is norm nonexpansive (operator norm $1$) but not norm-preserving unless $\beta_{t}\in\{0,2\}$. Singular values equal the absolute eigenvalues because $\mathbf{H}_{t}$ is symmetric; the component along $\mathbf{w}_{t}$ is scaled by $|1-\beta_{t}|<1$ for any $\beta_{t}\in(0,2)\setminus\{0,2\}$, and flips sign when $\beta_{t}>1$ (a design choice in PaTH to allow negative eigenvalues for state-tracking).
@@
Hence $\mathbf{P}_{j\to i}$ is (at best) nonexpansive, with a worst-case exponential lower bound on the smallest singular value governed by the path-length product of $|1-\beta_{s}|$. Whenever some $\beta_{s}$ is close to $1$, $\mathbf{H}_{s}$ is nearly singular (and exactly singular if $\beta_{s}=1$), driving $\sigma_{\min}(\mathbf{P}_{j\to i})$ toward zero. Volume contraction is quantified by
@@
Implications for long-context modeling. Because the PaTH transport multiplies the Q/K bilinear by $\mathbf{P}_{j\to i}$, any persistent deviation of $\beta_{t}$ from $\{0,2\}$ yields cumulative energy loss along a moving one-dimensional subspace. This concentrates mass in progressively fewer directions and can flatten or attenuate long-range logits $\mathbf{q}_{i}^{\top}\mathbf{P}_{j\to i}\mathbf{k}_{j}$ as iji-j grows, unless additional renormalizations or forget-gates are introduced. In contrast, GRAPE-M maps lie in $\operatorname{SO}(d)$, so for both non‑contextual and contextual types, all singular values are $1$; volumes and norms are preserved, and Lyapunov exponents are 0, avoiding contraction‑induced degradation of long‑range interactions.